# Generalization of DNNs:

### v1

@@ Appendix C Composition of Sobolev Balls
Given a distribution $\pi$ with support in $B(0,r)$, we have that with probability $1-\delta$ for all functions $f\in\mathcal{F}=\left\{f:\left\|f\right\|_{W^{\nu,2}}\leq R,\left\|f\right\|_{\infty}\leq R\right\}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)\leq 2C_{1}RE_{\nicefrac{{\nu}}{{d}}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
where $E_{r}(N)=N^{-\frac{1}{2}}$ if $r>\frac{1}{2}$, $E_{r}(N)=N^{-\frac{1}{2}}\log N$ if $r=\frac{1}{2}$, and $E_{r}(N)=N^{-r}$ if $r<\frac{1}{2}$.
@@
(1) We know from Theorem 5.2 of [9] that the Sobolev ball $B_{W^{\nu,2}}(0,R)$ over any $d$-dimensional hypercube $\Omega$ satisfies
$\log\mathcal{N}_{2}(B_{W^{\nu,2}}(0,R),\pi,\epsilon)\leq C_{0}\left(\frac{R}{\epsilon}\right)^{\frac{d}{\nu}}$
@@
(2) By Dudley’s theorem we can bound the Rademacher complexity of our function class $\mathcal{B}(X)$ evaluated on any training set $X$:
$R(\mathcal{B}(X))$ $\leq R2^{-M}+\frac{6R}{\sqrt{N}}\sum_{k=1}^{M}2^{-k}\sqrt{C_{0}\left(\frac{R}{R2^{-k}}\right)^{\frac{d}{\nu}}}$
$=R2^{-M}+\frac{6R}{\sqrt{N}}\sqrt{C_{0}}\sum_{k=1}^{M}2^{k(\frac{d}{2\nu}-1)}.$
If $2\nu=d$, we take $M=\frac{1}{2}\log N$ and obtain the bound
$\frac{R}{\sqrt{N}}+\frac{6R}{\sqrt{N}}\sqrt{C_{0}}\frac{1}{2}\log N\leq C_{1}R\frac{\log N}{\sqrt{N}}.$
If $2\nu>d$, we take $M=\infty$ and obtain the bound
$\frac{6R}{\sqrt{N}}\sqrt{C_{0}}\left(\frac{2^{\frac{d}{2\nu}-1}}{1-2^{\frac{d}{2\nu}-1}}\right)\leq C_{1}R\frac{1}{\sqrt{N}}.$
If $2\nu<d$, we take $M=\frac{\nu}{d}\log N$ and obtain the bound
$R2^{-M}+\frac{6R}{\sqrt{N}}\sqrt{C_{0}}2^{\frac{d}{2\nu}-1}\left(\frac{2^{M(\frac{d}{2\nu}-1)}-1}{2^{\frac{d}{2\nu}-1}-1}\right)\leq C_{1}RN^{-\frac{\nu}{d}}.$
Putting it all together, we obtain that $R(\mathcal{B}(X))\leq C_{1}E_{\nicefrac{{\nu}}{{d}}}(N)$.
(3) For any $\rho$-Lipschitz loss function $\ell(x,f(x))$ with $\left|\ell(x,y)\right|\leq c_{0}$, we know that with probability $1-\delta$ over the sampling of the training set $X$ from the distribution $\pi$, we have for all $f\in\mathcal{F}$
$\mathbb{E}_{x\sim\pi}\left[\ell(x,f(x))\right]-\frac{1}{N}\sum_{i=1}^{N}\ell(x_{i},f(x_{i}))$ $\leq 2\mathbb{E}_{X^{\prime}}\left[R(\ell\circ\mathcal{F}(X^{\prime}))\right]+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
$\leq 2C_{1}E_{\nicefrac{{\nu}}{{d}}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
@@
Let $\mathcal{F}=\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1}$ where =$\mathcal{F}_{\ell}=\left\{f_{\ell}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}\text{ s.t. }\left\|f_{\ell}\right\|_{W^{\nu_{\ell},2}}\leq R_{\ell},\left\|f_{\ell}\right\|_{\infty}\leq b_{\ell},Lip(f_{\ell})\leq\rho_{\ell}\right\}$, and let $r^{*}=\min_{\ell}r_{\ell}$ for $r_{\ell}=\frac{\nu_{\ell}}{d_{\ell-1}}$, then with probability $1-\delta$ we have for all $f\in\mathcal{F}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)$ $\leq\rho C_{0}\left(\sum_{\ell=1}^{L}\left(C_{\ell}\rho_{L:\ell+1}R_{\ell}\right)^{\frac{1}{r^{*}+1}}\right)^{r^{*}+1}E_{r^{*}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}},$
where $C_{\ell}$ depends only on $d_{\ell-1},d_{\ell},\nu_{\ell},b_{\ell-1}$.
@@
(1) We know from Theorem 5.2 of [9] that the Sobolev ball $B_{W^{\nu_{\ell},2}}(0,R_{\ell})$ over any $d_{\ell}$-dimensional hypercube $\Omega$ satisfies
$\log\mathcal{N}_{2}(B_{W^{\nu,2}}(0,R_{\ell}),\pi_{\ell-1},\epsilon_{\ell})\leq\left(C_{\ell}\frac{R_{\ell}}{\epsilon_{\ell}}\right)^{\frac{1}{r_{\ell}}}$
for a constant $C_{\ell}$ that depends on the size of hypercube and the dimension $d_{\ell}$ and the regularity $\nu_{\ell}$ and any measure $\pi_{\ell-1}$ supported in the hypercube.
@@
$\log\mathcal{N}_{2}(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1},\sum_{\ell=1}^{L}\rho_{L:\ell+1}\epsilon_{\ell})\leq\sum_{\ell=1}^{L}\left(C_{\ell}\frac{R_{\ell}}{\epsilon_{\ell}}\right)^{\frac{1}{r_{\ell}}}.$
Given $r^{*}=\min_{\ell}r_{\ell}$, we can bound it by $\sum_{\ell=1}^{L}\left(C_{\ell}\frac{R_{\ell}}{\epsilon_{\ell}}\right)^{\frac{1}{r^{*}}}$ and by then choosing $\epsilon_{\ell}=\frac{\rho_{L:\ell+1}^{-1}\left(\rho_{L:\ell+1}C_{\ell}R_{\ell}\right)^{\frac{1}{r^{*}+1}}}{\sum_{\ell}\left(\rho_{L:\ell+1}C_{\ell}R_{\ell}\right)^{\frac{1}{r^{*}+1}}}\epsilon$, we obtain that
$\log\mathcal{N}_{2}(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1},\epsilon)\leq\left(\sum_{\ell=1}^{L}\left(\rho_{L:\ell+1}C_{\ell}R_{\ell}\right)^{\frac{1}{r^{*}+1}}\right)^{r^{*}+1}\epsilon^{-\frac{1}{r^{*}}}.$
(2,3) It the follows by a similar argument as in points (2, 3) of the proof of Proposition 8 that there is a constant $C_{0}$ such that with probability $1-\delta$ for all $f\in\mathcal{F}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)$ $\leq C_{0}\left(\sum_{\ell=1}^{L}\left(\rho_{L:\ell+1}C_{\ell}R_{\ell}\right)^{\frac{1}{r^{*}+1}}\right)^{r^{*}+1}E_{r^{*}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
Appendix D Generalization at the Regularized Global Minimum
In this section, we first give the proof of Theorem 5 and then present detailed proofs of lemmas used in the proof. The lemmas are largely inspired by [5] and may be of independent interest.
D.1 Theorem 5 in Section 4.2
@@ Appendix D Generalization at the Regularized Global Minimum
Given a true function $f_{L^{*}:1}^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ going through the dimensions $d_{0}^{*},\dots,d_{L^{*}}^{*}$, along with a continuous input distribution $\pi_{0}$ supported in $B(0,b_{0})$, such that the distributions $\pi_{\ell}$ of $f_{\ell}^{*}(x)$ (for $x\sim\pi_{0}$) are continuous too and supported inside $B(0,b_{\ell})\subset\mathbb{R}^{d_{\ell}^{*}}$. Further assume that there are differentiabilities $\nu_{\ell}$ and radii $R_{\ell}$ such that $\left\|f_{\ell}^{*}\right\|_{W^{\nu_{\ell},2}(B(0,b_{\ell}))}\leq R_{\ell}$, and $\rho_{\ell}$ such that $Lip(f_{\ell}^{*})\leq\rho_{\ell}$. For an infinite width AccNet with $L\geq L^{*}$ and dimensions $d_{\ell}\geq d_{1}^{*},\dots,d_{L^{*}-1}^{*}$, we have for the ratios $\tilde{r}_{\ell}=\frac{\nu_{\ell}}{d_{\ell}^{*}+3}$:
(1) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda R(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=\tilde{O}(N^{-\min\{\frac{1}{2},\tilde{r}_{1},\dots,\tilde{r}_{L^{*}}\}})$.
(2) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda\tilde{R}(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=\tilde{O}(N^{-\frac{1}{2}+\sum_{\ell=1}^{L^{*}}\max\{0,\tilde{r}_{\ell}-\frac{1}{2}\}})$.
@@
Each $f_{\ell}^{*}$ can be approximated by another function $\tilde{f}_{\ell}$ with bounded $F_{1}$-norm and Lipschitz constant. Actually if $32\nu_{\ell}^{*}\geq d_{\ell-1}^{*}+$ 3 one can choose $\tilde{f}_{\ell}=f_{\ell}^{*}$ since $\left\|f_{\ell}^{*}\right\|_{F_{1}}\leq C_{\ell}R_{\ell}$ by Lemma 14, and by assumption $Lip(\tilde{f}_{\ell})\leq\rho_{\ell}$. If $32\nu_{\ell}^{*}<d_{\ell-1}^{*}+$ 3, then by Lemma 13 we know that there is a $\tilde{f}_{\ell}$ with $\left\|\tilde{f}_{\ell}\right\|_{F_{1}}\leq C_{\ell}R_{\ell}\epsilon_{\ell}^{-\frac{1}{2\tilde{r}_{\ell}}+1}$ and $Lip(\tilde{f}_{\ell})\leq C_{\ell}Lip(f_{\ell}^{*})\leq C_{\ell}\rho_{\ell}$ and error
@@
For any $L\geq L^{*}$, dimensions $d_{\ell}\geq d_{\ell}^{*}$ and widths $w_{\ell}\geq N$, we can build an AccNet that fits eactly $\tilde{f}_{L:1}$, by simply adding zero weights along the additional dimensions and widths, and by adding identity layers if $L>L^{*}$, since it is possible to represent the identity on $\mathbb{R}^{d}$ with a shallow network with $2d$ neurons and $F_{1}$-norm $2d$ (by having two neurons $e_{i}\sigma(e_{i}^{T}\cdot)$ and $-e_{i}\sigma(-e_{i}^{T}\cdot)$ for each basis $e_{i}$). Since the cost in parameter norm of representing the identity scales with the dimension, it is best to add those identity layers at the minimal dimension $\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}$. We therefore end up with a AccNet with $L-L^{*}$ identity layers (with $F_{1}$ norm $2\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}$) and $L^{*}$ layers that approximate each of the $f_{\ell}^{*}$ with a bounded $F_{1}$-norm function $\tilde{f}_{\ell}$.
Since $f_{L:1}^{*}$ has zero population loss, the population loss of the AccNet $\tilde{f}_{L:1}$ is bounded by $\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}$. By McDiarmid’s inequality, we know that with probability $1-\delta$ over the sampling of the training set, the training loss is bounded by $\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+B\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$. (1) The global minimizer $\hat{f}_{L:1}=\hat{f}_{L}\circ\cdots\circ\hat{f}_{1}$ of the regularized loss (with the first regularization term) is therefore bounded by
$\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+B\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
@@
$+\lambda\sqrt{2d}\prod_{\ell=1}^{L^{*}}C_{\ell}\rho_{\ell}\left[\sum_{\ell=1}^{L^{*}}\frac{1}{C_{\ell}\rho_{\ell}}\begin{cases}C_{\ell}R_{\ell}&2\nu_{\ell}^{*}\geq d_{\ell-1}^{*}+3\\ C_{\ell}R_{\ell}\epsilon_{\ell}^{-\frac{1}{2\tilde{r}_{\ell}}+1}&2\nu_{\ell}^{*}<d_{\ell-1}^{*}+3\end{cases}+2(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right].$
@@
Taking $\epsilon_{\ell}=E_{\tilde{r}_{min}}(N)$ and $\lambda=N^{-\frac{1}{2}}\log N$, this is upper bounded by
$\left[\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}+C\sqrt{2d}r\prod_{\ell=1}^{L^{*}}C_{\ell}\rho_{\ell}\sum_{\ell=1}^{L^{*}}\frac{R_{\ell}}{\rho_{\ell}}+2(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right]E_{\tilde{r}_{min}}(N)+B\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
@@
which implies that at the globla minimizer of the regularized loss, the (unregularized) train loss is of order $E_{\tilde{r}_{min}}(N)$ and the complexity measure $R(\hat{f}_{1},\dots,\hat{f}_{L})$ is of order $\frac{1}{N}E_{\tilde{r}_{min}}(N)$ which implies that the test error will be of order
$\mathcal{L}(f)$ $\leq\left[2\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}+2C\sqrt{2d}r\prod_{\ell=1}^{L^{*}}C_{\ell}\rho_{\ell}\sum_{\ell=1}^{L^{*}}\frac{R_{\ell}}{\rho_{\ell}}+2(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right]E_{\tilde{r}_{min}}(N)$
$+(2B+c_{0})\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
@@
$\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+B\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
$+2\lambda\sqrt{2d}\prod_{\ell=1}^{L^{*}}\begin{cases}C_{\ell}R_{\ell}&2\nu_{\ell}^{*}\geq d_{\ell-1}^{*}+3\\ C_{\ell}R_{\ell}\epsilon_{\ell}^{-\frac{1}{2\tilde{r}_{\ell}}+1}&2\nu_{\ell}^{*}<d_{\ell-1}^{*}+3\end{cases}\left[L^{*}+(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right],$

### v2

@@ Appendix D Composition of Sobolev Balls
Given a distribution $\pi$ with support in $B(0,b)$, we have that with probability $1-p$ for all functions $f\in\mathcal{F}=\left\{f:\left\|f\right\|_{W^{\nu,2}}\leq R,\left\|f\right\|_{\infty}\leq B\right\}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $=c_{0}\left(\frac{B^{2}R^{r}}{N}\right)^{\frac{1}{2+r}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}},$
where $r=\frac{d}{\nu}$. Therefore if ~$\sqrt{\tilde{\mathcal{L}}_{N}(f)}\leq c_{0}\left(\frac{B^{2}R^{r}}{N}\right)^{\frac{1}{2+r}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$, we have
$\mathcal{L}(f)$ $\leq 8c_{0}^{2}\left(\frac{B^{2}R^{r}}{N}\right)^{\frac{2}{2+r}}+8c_{1}^{2}B^{2}\frac{-\log\nicefrac{{p}}{{2}}}{N}.$
@@
(1) We know from Theorem 5.2 of (Birman & Solomjak, 1967) that the Sobolev ball $B_{W^{\nu,2}}(0,R)$ over any $d$-dimensional hypercube $\Omega$ satisfies
$\log\mathcal{N}_{2}(B_{W^{\nu,2}}(0,R),\epsilon)\leq C_{0}\left(\frac{R}{\epsilon}\right)^{\frac{d}{\nu}}$
@@
(2) We now apply Theorem 7 along the lines of Example 1 with $r=\frac{d}{\nu}$ and $c=C_{0}R^{r}$, to obtain that
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq 2\left(2B^{2}\frac{C_{0}R^{r}}{N}\right)^{\frac{1}{2+r}}+(\sqrt{2}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$
$=c_{0}\left(\frac{B^{2}R^{r}}{N}\right)^{\frac{1}{2+r}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}.$
The bound on the covering number of Sobolev balls, can then be used to bound compositions of Sobolev balls, thanks to the following general result:
@@
We can now put the two previous results together to obtain a generalization bound for composition of Sobolev balls:
@@
Let $\mathcal{F}=\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1}$ where =$\mathcal{F}_{\ell}=\left\{f_{\ell}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}\text{ s.t. }\left\|f_{\ell}\right\|_{W^{\nu_{\ell},2}}\leq R,\left\|f_{\ell}\right\|_{\infty}\leq B,Lip(f_{\ell})\leq\rho_{\ell}\right\}$, and let $r^{*}=\min_{\ell}r_{\ell}$ for $r_{\ell}=\frac{\nu_{\ell}}{d_{\ell-1}}$, then with probability $1-p$ we have for all $f\in\mathcal{F}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq c_{0}\left(\frac{B^{2}R^{r^{*}}}{N}\right)^{\frac{1}{2+r^{*}}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}},$
where $c_{0},c_{1}$ depend only on $r_{\ell},\rho_{\ell}$ and $L$. Thus if ~$\sqrt{\tilde{\mathcal{L}}_{N}(f)}\leq c_{0}\left(\frac{B^{2}R^{r^{*}}}{N}\right)^{\frac{1}{2+r^{*}}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}$, we have
$\mathcal{L}(f)$ $\leq 8c_{0}^{2}\left(\frac{B^{2}R^{r^{*}}}{N}\right)^{\frac{2}{2+r^{*}}}+8c_{1}^{2}B^{2}\frac{-\log\nicefrac{{p}}{{2}}}{N}.$
@@
(1) We know from Theorem 5.2 of (Birman & Solomjak, 1967) that the Sobolev ball $B_{W^{\nu_{\ell},2}}(0,R)$ over any $d_{\ell}$-dimensional hypercube $\Omega$ satisfies
$\log\mathcal{N}_{2}(B_{W^{\nu,2}}(0,R),\pi_{\ell-1},\epsilon_{\ell})\leq\left(C_{\ell}\frac{R}{\epsilon_{\ell}}\right)^{r_{\ell}}$
for $r_{\ell}=\frac{d_{\ell}}{\nu_{\ell}}$ and a constant $C_{\ell}$ that depends on the size of hypercube and the dimension $d_{\ell}$ and the regularity $\nu_{\ell}$ and any measure $\pi_{\ell-1}$ supported in the hypercube.
@@
$\log\mathcal{N}_{2}(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1},\sum_{\ell=1}^{L}\rho_{L:\ell+1}\epsilon_{\ell})\leq\sum_{\ell=1}^{L}\left(C_{\ell}\frac{R}{\epsilon_{\ell}}\right)^{r_{\ell}}.$
Given $r^{*}=\max_{\ell}r_{\ell}$, we can bound it by $R^{r^{*}}\sum_{\ell=1}^{L}\left(C_{\ell}\frac{1}{\epsilon_{\ell}}\right)^{r^{*}}$ and by then choosing $\epsilon_{\ell}=\frac{\rho_{L:\ell+1}^{-1}\left(\rho_{L:\ell+1}C_{\ell}\right)^{\frac{r^{*}}{1+r^{*}}}}{\sum_{\ell}\left(\rho_{L:\ell+1}C_{\ell}\right)^{\frac{r^{*}}{1+r^{*}}}}\epsilon$, we obtain that
$\log\mathcal{N}_{2}(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1},\epsilon)\leq\left(\sum_{\ell=1}^{L}\left(\rho_{L:\ell+1}C_{\ell}\right)^{\frac{r^{*}}{1+r^{*}}}\right)^{1+r^{*}}\left(\frac{R}{\epsilon}\right)^{r^{*}}.$
(2) Again, we apply Theorem 7 along the lines of Example 1, we obtain that
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq c_{0}\left(\frac{B^{2}R^{r^{*}}}{N}\right)^{\frac{1}{2+r^{*}}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}.$
Appendix E Generalization at the Regularized Global Minimum
In this section, we first give the proof of Theorem 5 and then present detailed proofs of lemmas used in the proof. The lemmas are largely inspired by (Bach, 2017) and may be of independent interest.
E.1 Theorem 5 in Section 4.2
@@ Appendix E Generalization at the Regularized Global Minimum
Given a true function $f_{L^{*}:1}^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ going through the dimensions $d_{0}^{*},\dots,d_{L^{*}}^{*}$, along with a continuous input distribution $\pi_{0}$ supported in $B(0,b_{0})$, such that the distributions $\pi_{\ell}$ of $f_{\ell}^{*}(x)$ (for $x\sim\pi_{0}$) are continuous too and supported inside $B(0,b_{\ell})\subset\mathbb{R}^{d_{\ell}^{*}}$. Further assume that there are differentiabilities $\nu_{\ell}$ and radii $R_{\ell}$ such that $\left\|f_{\ell}^{*}\right\|_{W^{\nu_{\ell},2}(B(0,b_{\ell}))}\leq R_{\ell}$, and $\rho_{\ell}$ such that $Lip(f_{\ell}^{*})\leq\rho_{\ell}$. For an infinite width AccNet with $L\geq L^{*}$ and dimensions $d_{\ell}\geq d_{1}^{*},\dots,d_{L^{*}-1}^{*}$, we have for the ratios $\tilde{r}_{\ell}=\max\{\frac{d_{\ell-1}^{*}+3}{\nu_{\ell}},2\}$:
(1) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda R(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=O(N^{-\frac{2}{2+\tilde{r}_{\max}}})$ for $\tilde{r}_{\max}=\max\{\tilde{r}_{1},\dots,\tilde{r}_{L}\}$.
(2) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda\tilde{R}(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=O(N^{-\frac{2}{2+\tilde{r}_{sum}}})$ for $\tilde{r}_{sum}=2+\sum_{\ell}\tilde{r}_{\ell}-2$.
@@
Each $f_{\ell}^{*}$ can be approximated by another function $\tilde{f}_{\ell}$ with bounded $F_{1}$-norm and Lipschitz constant. Actually if $32\nu_{\ell}\geq d_{\ell-1}^{*}+$ 3 one can choose $\tilde{f}_{\ell}=f_{\ell}^{*}$ since $\left\|f_{\ell}^{*}\right\|_{F_{1}}\leq C_{\ell}R_{\ell}$ by Lemma 16, and by assumption $Lip(\tilde{f}_{\ell})\leq\rho_{\ell}$. If $32\nu_{\ell}<d_{\ell-1}^{*}+$ 3, then by Lemma 15 we know that there is a $\tilde{f}_{\ell}$ with $\left\|\tilde{f}_{\ell}\right\|_{F_{1}}\leq C_{\ell}R_{\ell}\epsilon_{\ell}^{-\frac{d^{*}_{\ell-1}+3}{2\nu_{\ell}}+1}$ and $Lip(\tilde{f}_{\ell})\leq C_{\ell}Lip(f_{\ell}^{*})\leq C_{\ell}\rho_{\ell}$ and error
@@
Note that by setting $\tilde{r}_{\ell}=\max\{\frac{d_{\ell-1}^{*}+3}{\nu_{\ell}},2\}$, we can write $\left\|\tilde{f}_{\ell}\right\|_{F_{1}}\leq C_{\ell}R_{\ell}\epsilon_{\ell}^{-\frac{\tilde{r}_{\ell}}{2}+1}$ in both cases.
@@
For any $L\geq L^{*}$, dimensions $d_{\ell}\geq d_{\ell}^{*}$ and widths $w_{\ell}\geq N$, we can build an AccNet that fits exactly $\tilde{f}_{L:1}$, by simply adding zero weights along the additional dimensions and widths, and by adding identity layers if $L>L^{*}$, since it is possible to represent the identity on $\mathbb{R}^{d}$ with a shallow network with $2d$ neurons and $F_{1}$-norm $2d$ (by having two neurons $e_{i}\sigma(e_{i}^{T}\cdot)$ and $-e_{i}\sigma(-e_{i}^{T}\cdot)$ for each basis $e_{i}$). Since the cost in parameter norm of representing the identity scales with the dimension, it is best to add those identity layers at the minimal dimension $\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}$. We therefore end up with a AccNet with $L-L^{*}$ identity layers (with $F_{1}$ norm $2\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}$) and $L^{*}$ layers that approximate each of the $f_{\ell}^{*}$ with a bounded $F_{1}$-norm function $\tilde{f}_{\ell}$.
Since $f_{L:1}^{*}$ has zero population loss, the population loss of the AccNet $\tilde{f}_{L:1}$ is bounded by $(\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell})^{2}$.Using Bennett’s inequality as in the proof of Theorem 7 (though in the ‘other direction’ since we want to bound ~$\tilde{\mathcal{L}}_{N}$ in terms of $\mathcal{L}$ and not vice versa), we obtain that with probability $1-p$
~$\tilde{\mathcal{L}}_{N}(\tilde{f}_{L:1})$ $\leq\mathcal{L}(\tilde{f}_{L:1})+\sqrt{8B^{2}\mathcal{L}(\tilde{f}_{L:1})\frac{-\log p}{N}}+8B^{2}\frac{-\log p}{N}$
$\leq\left(\sqrt{\mathcal{L}(\tilde{f}_{L:1})}+\sqrt{8B^{2}\frac{-\log p}{N}}\right)^{2}$
$\leq\left(\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+\sqrt{8B^{2}\frac{-\log p}{N}}\right)^{2}$
(1) At the global minimizer $\hat{f}_{L:1}=\hat{f}_{L}\circ\cdots\circ\hat{f}_{1}$ of the regularized loss (with the first regularization term), the regularized loss is upper bounded by
$\left(\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+\sqrt{8B^{2}\frac{-\log p}{N}}\right)^{2}$
@@
$+\lambda\sqrt{2d}\prod_{\ell=1}^{L^{*}}C_{\ell}\rho_{\ell}\left[\sum_{\ell=1}^{L^{*}}\frac{R_{\ell}\epsilon_{\ell}^{\frac{2-\tilde{r}_{\ell}}{2}}}{C_{\ell}\rho_{\ell}}+2(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right].$
@@
Taking $\epsilon_{\ell}=N^{-\frac{1}{2+\tilde{r}_{\ell}}}$ and $\lambda=N^{-\frac{1}{2}}$, this is upper bounded by
$\left(\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}N^{-\frac{1}{2+\tilde{r}_{\ell}}}+\sqrt{8B^{2}\frac{-\log p}{N}}\right)^{2}$
$+\sqrt{2d}\prod_{\ell=1}^{L^{*}}C_{\ell}\rho_{\ell}\left[\sum_{\ell=1}^{L^{*}}\frac{R_{\ell}}{C_{\ell}\rho_{\ell}}N^{-\frac{2}{2+\tilde{r}_{\ell}}}+2(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}N^{-\frac{1}{2}}\right].$
@@
The above is of order $N^{-\frac{2}{2+\tilde{r}_{\max}}}$ for $\tilde{r}_{\max}=\max\{\tilde{r}_{\ell}:\ell=1,\dots,L\}$, which implies that at the global minimizer of the regularized loss, the (unregularized) train loss is of order $N^{-\frac{2}{2+\tilde{r}_{\max}}}$ and the complexity measure $R(\hat{f}_{1},\dots,\hat{f}_{L})$ is of order $N^{\frac{1}{2}-\frac{2}{2+\tilde{r}_{\max}}}$ which implies that the test error will be of order $N^{-\frac{2}{2+\tilde{r}_{\max}}}$ as well.
@@
$\left(\rho\sum_{\ell=1}^{L}\rho_{L:\ell+1}C_{L:\ell+1}c_{\ell}\epsilon_{\ell}+\sqrt{8B^{2}\frac{-\log p}{N}}\right)^{2}$
$+\lambda\sqrt{2d}\left(L^{*}+(L-L^{*})\min\{d_{0}^{*},\dots,d_{L^{*}}^{*}\}\right)\prod_{\ell=1}^{L^{*}}C_{\ell}R_{\ell}\epsilon_{\ell}^{\frac{2-\tilde{r}_{\ell}}{2}}.$